- Press Release
- Dec 8, 2022
NASA International Space Station On-Orbit Status 17 August 2016
Flight controllers and the Expedition 48 crew are preparing for tonight’s International Docking Adapter extraction work and Friday morning’s installation spacewalk. The orbital residents are also continuing to load the SpaceX Dragon with gear and science for return to Earth.
Controllers on the ground checked the Canadian robotics systems they will use to remove the International Docking Adapter from the rear of the SpaceX Dragon tonight. The new adapter will be extracted with the Canadarm2 then maneuvered to a point about three feet away from its installation point. It will then be installed on the Pressurized Mating Adapter-2 which is attached to the forward end of the Harmony module.
The final and intricate installation work will be done during a 6.5 hour spacewalk scheduled to begin Friday at 8:05 a.m. EDT with astronauts Jeff Williams and Kate Rubins . The new adapter, the first of two, will enable new Commercial Crew vehicles being developed by Boeing and SpaceX to dock at the International Space Station in the future.
In the meantime, the SpaceX Dragon cargo craft is still being loaded with gear ahead of its return to Earth on Aug. 26. Rubins spent the morning packing Dragon with research samples and used hardware for analysis back on Earth.
On-Orbit Status Report
Fluid Shifts Operations In the Service Module: With ground team assistance, crewmembers performed Fluid Shifts Imaging exams by configuring the Optical Coherence Tomography (OCT) and the Distortion Product Otoacoustic Emission (DPOAE) hardware, before completing a DPOAE test, OCT exam, and a Tonometry exam. The purpose of this investigation is to characterize the space flight-induced fluid shift, including intra- and extravascular shifts, intra- and extracellular shifts, changes in total body water and lower vs. upper body shifts. Results from this investigation are expected to help define the causes of the ocular structure and vision changes associated with long duration space flight, and assist in the development of countermeasures.
Biological Research in Canisters Natural Products (BRIC-NP) Cold Stowage Preparation: The crew retrieved the BRIC-NP canisters from EXPRESS Rack 2 and inserted them in the Glacier for return on SpX-9. The BRIC-NP investigation, radiation-tolerant fungal strains isolated from the Chernobyl nuclear power plant are exposed to spaceflight conditions on board the ISS, then screened for the biological production of beneficial medical or agricultural substances.
eValuatIon And monitoring of microBiofiLms insidE the ISS (ViABLE) Payload Return: The crew photographed all four ViABLE Bags inside the Functional Cargo Block (FGB) locker before removing and inserting each one into separate ViABLE Return Bags and placing them into Ziploc bags for return on SpX-9. ViABLE involves the evaluation of microbial biofilm development on metallic and textile space materials located inside and on the cover of Nomex pouches. Microbial biofilms are known for causing damage and contamination on the Mir space station and the ISS. The potential application of novel methodologies and products to treat space materials may lead to improvements in the environmental quality of confined human habitats in space and on earth.
Habitability Human Factors Directed Observations: The crew recorded and submitted a walk-through video documenting observations of life onboard ISS, providing insight related to human factors and habitability. The Habitability investigation collects observations about the relationship between crew members and their environment on the ISS. Observations can help spacecraft designers understand how much habitable volume is required, and whether a mission’s duration impacts how much space crew members need.
Dose Tracker: The Dose Tracker app was configured and the crew completed entries for medication tracking on an iPad. This investigation documents the medication usage of crewmembers before and during their missions by capturing data regarding medication use during spaceflight, including side effect qualities, frequencies and severities. The data is expected to either support or counter anecdotal evidence of medication ineffectiveness during flight and unusual side effects experienced during flight. It is also expected that specific, near-real-time questioning about symptom relief and side effects will provide the data required to establish whether spaceflight-associated alterations in pharmacokinetics (PK) or pharmacodynamics (PD) is occurring during missions.
Oxygen Generation System (OGS) Hydrogen (H2) Sensor Remove & Replace (R&R): The crew completed OGS H2 sensor Orbital Replacement Unit (ORU) purge adapter operations, R&R of the H2 sensor ORU and AAA cleaning with inlet inspection and cleaning. This activity was scheduled due to ORU end-of-life.
Urine Processing Assembly (UPA) Sample Collection: The UPA Distillate Filter and Purge Filter were removed and replaced for return samples in support of the UPA elevated conductivity investigation. The preceding UPA process cycle had elevated conductivity levels and the samples will help ground teams understand the difference in conductivity with the purge line reconnected.
Waste and Hygiene Compartment (WHC) Check Separator Light: To isolate the cause of Check Separator Light illumination, the crew performed an inspection of the hoses and electrical connections on the Urine Receptacle. They reported that the electrical cable on the Urine Receptacle was coiled tightly and may be kinked. They removed the zip tie that was holding the cable together to allow more slack and demated/remated the XT2 connector. Additional troubleshooting steps are in work for the crew to perform as a result of yesterday’s Flight Investigation Team (FIT) recommendations.
Extravehicular Robotics Operations: Yesterday afternoon, Robotics Ground Controllers maneuvered the Space Station Remote Manipulator System (SSRMS) to unstow the Special Purpose Dexterous Manipulator (SPDM). Following SPDM unstow, they maneuvered the SSRMS to position it over the SpX-9 Dragon Trunk and configured the SPDM to extract the International Docking Adapter 2 (IDA2) from the trunk later today. SPDM checkouts were also completed in preparation for the IDA2 Extravehicular Activity (EVA) this Friday.
Today’s Planned Activities
All activities were completed unless otherwise noted.
Calf Volume Measurement / r/g 3080
FLUID SHIFTS. Comm configuration for the experiment / r/g 9995
FLUID SHIFTS. Gathering and Connecting Equipment for TV coverage
Soyuz 731 Samsung Tablet Recharge, initiate
Biological Research in Canisters Natural Products Stowage Preparation
Study of veins in lower extremities / r/g 3081
Microgravity Science Glovebox (MSG) Video file transfer to flash drive
BRICNP. Sample Insertion into Glacier
OGA Hydrogen Sensor R&R
Collecting surface samples from FGB equipment and structures / r/g 3083
Oxygen Generation System (OGS) Hydrogen Sensor ORU Purge Adapter (HOPA) Operations
FLUID SHIFTS. OCT Hardware Setup in SM
FLUID SHIFTS. Connecting [OCT] Laptop to BRI r/g 3104
DOSETRK iPad data entry
RSS2 Laptop SW Upgrade for Auto Data Downlink via RSPI r/g 3105
FLUID SHIFTS. OCT Power up in SM
Printing Housekeeping Procedure
DOSETRK Questionnaire Completion
Soyuz 720 Samsung Tablet Recharge, Initiate
Soyuz 731 Samsung Tablet Recharge, – Terminate
FLUID SHIFTS. DPOAE Setup in SM
FLUID SHIFTS. Operator Assistance with Chibis and Gamma-1 r/g 3103
FLUID SHIFTS. Chibis Setup / r/g 3103
FLUID SHIFTS. TONO Hardware setup in SM
OGA Hydrogen Sensor R&R, Part 2
FLUID SHIFTS. Data Gathering in the SM, Subject
FLUID SHIFTS. Gathering Data in SM, Operator
Installation of ??1? and Replacement of ??1? ??-9??-8 in FGB, Prep for Work, Clearing FGB panels r/g 3095
Fluid Shifts DPOAE Data Recovery
FLUID SHIFTS. Chibis Closeout Ops / r/g 3103
Water Recovery System Waste Water Tank Drain, Initiate
IMS Delta File Prep
IFM IMV Cleaning
FLUID SHIFTS. TONO SM Stowage
FLUID SHIFTS. OCT Power off in SM
OGA Hydrogen Sensor Cleaning
FLUID SHIFTS. Disconnecting OCT Laptop r/g 3106
DRAGON. Cargo Transfer Tagup
Water Recovery System (WRS) Drain Waste Water Tank, terminate
FLUID SHIFTS. ???? PDU De-Route / r/g 3106
FLUID SHIFTS. OCT Hardware Stowage
Study of veins in lower extremities / r/g 3081
Soyuz 720 Samsung tablet recharge, terminate
FLUID SHIFTS. CCFP/PC Stowage
Private Medical Conference
FLUID SHIFTS. Crew Onboard Support System (????) Hardware Deactivation and Closing Applications on CP SSC
MSG Transfer of Flash Drive File to Payload NAS
FLUID SHIFTS. Restoring nominal comm config
FLUID SHIFTS. Hardware Gathering before Transfer to USOS
MRM2 comm config to support the P/L Ops / Communications System
KULONOVSKIY KRISTALL. Experiment Session r/g 3102
FLUID SHIFTS. Hardware Transfer to USOS
UPA Gathering Brine Filter
FLUID SHIFTS. Cable removal
BIMS. Experiment Ops r/g 3094
BIMS. Operator Assistance During the Experiment / r/g 3094
FLUID SHIFTS. OCT and Laptop Setup
Monthly Inspection of T2 Treadmill System
Urine Processor Assembly (UPA) Brine Filter Changeout
MRM2 Comm Reconfig for Nominal Ops / Communications System
FLUID SHIFTS. Ultrasound 2 Setup and Activation
KULONOVSKIY KRISTALL. Copy and Downlink Data / r/g 3102
Robotic Work Station (RWS), Display and Control Panel (DCP) Checkout
FLUID SHIFTS. Ultrasound 2 Data Export
Crew Onboard Training (OBT), Self-Assessment Questionnaire
FLUID SHIFTS. Hardware stowage
HABIT Data Recording to iPad
Physical Fitness Evaluation (on the treadmill) r/g 3082
WRS Recycle Tank Fill from EDV
VIABLE. Bag preparation for return and photography
Countermeasures System (CMS), Sprint Exercise, Optional
On MCC Go Closeout ops with Komparus
Gathering Tools for IDA Installation
Relocating PBAs for upcoming EVA
Video Footage of Greetings / r/g 3107
Total Organic Carbon Analyzer (TOCA) Buffer Container (BC) Changeout
Completed Task List Items
All activities were completed unless otherwise noted.
ORU sensor R&R support
Nominal ground commanding
Three-Day Look Ahead:
Thursday, 08/18: EVA procedures review/tool config/Equipment Lock prep, Mouse Epigenetics cage maintenance
Friday, 08/19: IDA2 EVA
Saturday, 08/20: Post EVA cleanup activities, EVA debrief, Heart Cells media change, Mouse cage maintenance, CMO OBT
QUICK ISS Status – Environmental Control Group:
Component – Status
Elektron – On
Vozdukh – Manual
[???] 1 – SM Air Conditioner System (“SKV1”) – On
[???] 2 – SM Air Conditioner System (“SKV2”) – Off
Carbon Dioxide Removal Assembly (CDRA) Lab – Operate
Carbon Dioxide Removal Assembly (CDRA) Node 3 – Operate
Major Constituent Analyzer (MCA) Lab – Idle
Major Constituent Analyzer (MCA) Node 3 – Operate
Oxygen Generation Assembly (OGA) – Process
Urine Processing Assembly (UPA) – Norm
Trace Contaminant Control System (TCCS) Lab – Off
Trace Contaminant Control System (TCCS) Node 3 – Full Up